### 27 Jan 2024 v1

# Polynomial time auditing of statistical subgroup fairness for Gaussian data

Abstract
We study the problem of auditing classifiers with the notion of statistical subgroup fairness. [1] has shown that the problem of auditing combinatorial subgroups fairness is as hard as agnostic learning. Essentially all work on remedying statistical measures of discrimination against subgroups assumes access to an oracle for this problem, despite the fact that no efficient algorithms are known for it. If we assume the data distribution is Gaussian, or even merely log-concave, then a recent line of work has discovered efficient agnostic learning algorithms for halfspaces. Unfortunately, the boosting-style reductions given by Kearns et al. required the agnostic learning algorithm to succeed on reweighted distributions that may not be log-concave, even if the original data distribution was. In this work, we give positive and negative results on auditing for the Gaussian distribution: On the positive side, we an alternative approach to leverage these advances in agnostic learning and thereby obtain the first polynomial-time approximation scheme (PTAS) for auditing nontrivial combinatorial subgroup fairness: we show how to audit statistical notions of fairness over homogeneous halfspace subgroups when the features are Gaussian. On the negative side, we find that under cryptographic assumptions, no polynomial-time algorithm can guarantee any nontrivial auditing, even under Gaussian feature distributions, for general halfspace subgroups.
1 INTRODUCTION
The deployment of decision rules obtained using machine learning has raised the risk that the rules may exhibit biases against historically marginalized communities. In particular, [1] raised the concern that these decision rules may be biased against sub-groups characterized by a combination of “protected” attributes. Since there is an exponential number of such subgroups, even detecting such statistical patterns of discrimination is a nontrivial computational problem; indeed, [1] showed that the problem of finding disadvantaged subgroups is equivalent to the problem of agnostic learning, which is believed to be intractable in general for albeit the simplest classes of sets. Essentially all work [1, 2, 3] on remedying statistical measures of discrimination against subgroups assumes access to an oracle for this problem, despite the fact that no efficient algorithms are known for it. In this work we are proposing a solution for the fairness auditing problem with provable guarantees of efficiency and correctness.
1.1 Background and Motivation
Fairness learning has received massive attention in recent years. It turns out learning a fair classifier, in most cases, is equivalent to auditing [1, 2, 3]. In particular, if auditing is possible, learning a fair classifier is easy. There are many successful examples of fairness learning with auditing over a relatively small number of predetermined subgroups [4, 5]. However, a small number of predetermined subgroups, in many cases, is not enough to cover all the natural subgroups.
Example 1.1.
In the court case “DeGraffenreid v General Motors" [6], five Black women brought suit against General Motors for its discrimination against the group of Black women. Although no sex discrimination was revealed, the evidence showed that Black women hired after 1970 were discriminated against by the company’s seniority system. Such discrimination can be better demonstrated by an example shown in table 1. In particular, the hiring rate of a company could seemingly be fair in terms of gender of race alone but clearly discriminates against the subgroups of white men and black women. As a result, the court rejected the plaintiffs’ attempt to bring a suit not on behalf of Blacks or women, but specifically on behalf of Black women. In the ruling, the judge was specifically concerned about the proliferation of protected classes.
Table 1: an example of discrimination against subgroups
men women total
black 50 0 50
white 0 50 50
total 50 50 100
More generally, a classifier may appear to be fair on each individual attribute, e.g., gender, race, age, incomes, etc., and yet perform unfairly on subgroups defined on multiple attributes, i.e., the conjunction of such attributes. In the case of DeGraffenreid v General Motors, it is the conjunction of race and gender being discriminated against. The possible number of the conjunctions grows exponentially as the number of the “protected” attributes increases.
Thereafter, [1] proposed more general notions of statistical fairness that require auditing over subgroups defined on simple combinations of data features. Specifically, such combinations of features can be any simple representations, such as conjunctions and halfspaces, which, however, can generate exponentially many subgroups. They also showed that the problem of auditing subgroups defined by such simple representation is as hard as “weak agnostic learning” [7, 8]. While the problem of weak agnostic learning is widely believed to be computationally intractable [8, 9], its hardness does not necessarily hold for specific distributions. Thus, it is natural for someone to think of attacking the auditing problem through distribution-specific agnostic learning approaches as agnostic learning is a much more extensively studied problem. However, it turns out there are still obstacles remaining for doing so.
1.2 Challenges of Auditing through Agnostic Learning
The main challenge that prevents us from applying existing agnostic learning techniques to perform auditing based on the reduction by [1] is that it is formulated in terms of weak agnostic learning, that is, finding classifiers with error rates that are nonnegligibly better than guessing, and correspondingly weak auditing guarantees. What we would like instead is for the unfairness certificate to be near-optimal, so that the unfairness it exhibits certifies a bound on the unfairness of the classifier for all subgroups. When we have guarantees for arbitrary distributions, “boosting” [10] enables us to obtain high accuracy from such weak learners. Unfortunately, these techniques require re-weighting the data examples after which the distribution-specific properties may no longer hold.
One might hope to dodge this issue by casting the problem of finding a harmed subgroup as a Mixed-Integer Program and using solvers that, though they lack polynomial-time guarantees, obtain adequate performance in practice. In such an approach, the failure of the solver to find a feasible solution to the optimization problem is taken as the proof that the classifier is fair. Unfortunately, these solvers owe their speed in part to a lack of soundness, both due to numerical issues [11] and the complexity of the heuristics used to prune the search [12, 13], and it remains an open question whether or not it is possible to obtain acceptable performance (using the various advanced techniques employed by commercial solvers) while retaining the guarantee that the solver correctly reports infeasibility [14]. In any case, the works by [1, 15] and [16] that empirically studied these approaches to obtaining fair classifiers used linear regression as a proxy for the agnostic learning or cost-sensitive classification subroutines. Unfortunately, these heuristics do not even provide in-principle guarantees.
In this paper, we will present an alternative auditing approach for subgroups determined by homogeneous halfspaces with provable guarantees.
1.3 Our Contribution
Our main contribution is a more careful analysis of the relationship between auditing and agnostic learning: Given a fixed positive classification rate, the harm (w.r.t. statistical parity) suffered by a subgroup is affinely related to the error rate of the subgroup indicator. Thus, a solution to the agnostic learning problem directly gives a harmed subgroup. It is worth noting that under a standard normal distribution, the subclass of halfspaces with a fixed positive classification rate is given by the halfspaces with unit normal vectors and the same threshold.
Remark 1.2.
Our reduction to learning halfspaces with fixed positive classification rates can achieve arbitrarily high precision auditing and does not rely on re-weighting data examples or make any assumptions on the potentially unfair classifiers. This enables the use of the existing distribution-specific agnostic learning methods for auditing.
Furthermore, following [17], we also obtain a lower bound on the unfairness detectable when auditing for halfspace subgroups under Gaussian distributions by reducing the problem of continuous Learning With Errors (cLWE) to auditing.
For our algorithmic results, we will present a general auditing framework given an oracle for (distribution-specific) agnostic learning. Also, we give a randomized PTAS auditing algorithm for subgroups determined by homogeneous halfspaces under Gaussian data by applying the method from [18].
Remark 1.3.
We stress that a PTAS for auditing subgroups defined by homogeneous halfspaces for Gaussian distributions is, in fact, the best guarantee we know so far, hence, not trivial.
At first blush, the reliance on a (prima facie unverifiable) distributional assumption for the analysis of our auditing algorithm may seem to be at odds with our desire to certify the fairness of a classifier. Nevertheless, a line of recent works by [19] and [20] have shown that the properties of the data that are crucial to these algorithms for distribution-specific learning of halfspaces can be verified. Thus, these methods give a way of certifying fairness for families of nice distributions: so long as the data passes these tests and the audit reveals no subgroup that is significantly harmed, we may guarantee that the classifier is fair.
This paper will be organized as follow. Some necessary notations for our arguments are given in section 2. Then, we will present the main reduction and the hardness result in section 3. Section 4 will present out auditing framework as well as the distribution-specific PTAS algorithm. At last, we will discuss the limitation of our approach and our future work.
1.4 Related Work
[1] gave an game theory based approach for learning representations with subgroup fairness by assuming there exists an efficient oracle for auditing. A follow-up study [15] was later conducted to estimate their algorithm on real-world datasets. [3] showed a method of obtaining “multi-accurate” representations by assuming the existence of an efficient auditing oracle. Besides, [2] proposed a variant of statistical fairness called “multi-fairness,” which allows them to efficiently learn a multi-fair classifier with querying “relative fairness” of data pairs. As we discussed previously, the auditing oracles in these works were provided by using linear regression as a heuristic for the optimal halfspace, which does not provide guarantees. They also did not consider auditing for specific families of distributions. On the other hand, the works on agnostic learning for specific families of distributions, e.g., [21, 22, 18, 23, 24] do not consider how their techniques may be applied to the subgroup fairness auditing problem.
2 PRELIMINARIES
We use lowercase bold font characters to represent real vectors and subscripts to index the coordinates of each vector, e.g., $\mathbf{x}_{i}$ represents the $i$-th coordinate of vector $\mathbf{x}$. We denote the $l_{p}$-norm by $\lVert{\mathbf{x}}\rVert_{p}=\left(\sum_{i}\mathbf{x}_{i}^{p}\right)^{1/p}$, and $\bar{\mathbf{x}}=\mathbf{x}/\lVert{\mathbf{x}}\rVert_{2}$. We model each individual as a vector of protected attributes, i.e., $\mathbf{x}\in\mathcal{X}$.
Further, the probability of an event under a distribution $\mathcal{D}$ is denoted by $\Pr_{\mathbf{x}\sim\mathcal{D}}\{\cdot\}$. Specifically, $\mathcal{N}(0,\mathrm{I})$ denotes a standard normal distribution, where $\mathrm{I}$ represents the identity matrix. For simplicity of notation, we may use $\mathcal{N},\mathcal{N}_{\sigma}$ instead of $\mathcal{N}(0,\mathrm{I}),\mathcal{N}(0,\sigma^{2}\mathrm{I})$ or even drop $\mathcal{D}$ and $\mathcal{N}$ from the subscript when it is clear from the context.
Fact 2.1 (Rotational Invariance).
For any real vector $\mathbf{u}$, if $\mathbf{x}\sim\mathcal{N}(0,\mathrm{I})$, then $\bar{\mathbf{u}}^{\top}\mathbf{x}\sim\mathcal{N}(0,1)$.
To understand the problem of fairness auditing, it is necessary to define fairness or unfairness precisely. In this work, we focus on the notion of Statistical Parity Subgroup Fairness (SPSF). Formally, we have the following definition.
Definition 2.2 (Statistical Parity Subgroup Fairness).
Fix any binary classifier $c\in\mathcal{C}$ such that $c:\mathbb{R}^{d}\rightarrow\{-1,+1\}$, data distribution $\mathcal{D}$, collection of subgroups $\mathcal{G}$, and parameter $\gamma\in[0,1]$. Define
$d_{\mathcal{D}}(c,g)=\Pr_{\mathbf{x}\sim\mathcal{D}}\{c(\mathbf{x})=1\}-\Pr_{\mathbf{x}\sim\mathcal{D}}\{c(\mathbf{x})=1\ |\ \mathbf{x}\in g\}$ (1)
We say that g does not satisfy $\gamma$-statistical parity fairness (or is $\gamma$-unfair) with respect to $\mathcal{D}$ and $\mathcal{G}$, if $\exists g\in\mathcal{G}$ such that
$\Pr_{\mathbf{x}\sim\mathcal{D}}\{\mathbf{x}\in g\}\left|d_{\mathcal{D}}(c,g)\right|\geq\gamma$ (2)
Equation (1) is a straightforward way to quantify how much the positive classification rate within a subgroup deviates from that of the overall population. The goal of fairness auditing is to develop an “auditing algorithm" to efficiently find such a certificate $g\in\mathcal{G}$ for any $c\in\mathcal{C}$ with sample access to $\mathcal{D}$.
Definition 2.3 ([1] Auditing Algorithm).
Under the notion of Statistical Parity Subgroup Fairness, fix a collection of group indicators $\mathcal{G}$ over the protected features, and any $\delta,\gamma,\gamma^{\prime}\in(0,1)$ such that $\gamma^{\prime}\leq\gamma$. A $(\gamma,\gamma^{\prime})$-auditing algorithm for $\mathcal{G}$ with respect to distribution $\mathcal{D}$ is an algorithm $\mathcal{A}$ such that for any classifier $h$, when given access the joint distribution $(\mathcal{D},h(\mathcal{D}))$, $\mathcal{A}$ runs in time $\mathrm{poly}(1/\gamma^{\prime},\log(1/\delta))$, and with probability $1-\delta$, outputs a $\gamma^{\prime}$-unfair certificate for $h$ whenever $h$ is $\gamma$-unfair with respect to $\mathcal{D}$ and $\mathcal{G}$. If $h$ is $\gamma^{\prime}$-fair, $\mathcal{A}$ will output “fair".
Since our reduction involves the subclass of halfspaces with a “probabilistic lock”, we give the formal definition of it as follows.
Definition 2.4 (Halfspaces With Probabilistic Lock).
We use $\mathcal{H}^{d}$ to represent the collection of all the halfspaces in $\mathbb{R}^{d}$. Then, for any arbitrary distribution $\mathcal{D}$ over $\mathbb{R}^{d}$, we define the collection of all halfspaces with the same probability threshold $\mu$ as
$\mathcal{H}_{\mu}^{\mathcal{D}}:=\{h\in\mathcal{H}^{d}\ |\ \Pr_{\mathbf{x}\in\mathcal{D}}\{h(\mathbf{x})=1\}=\mu\}$ (3)
For conciseness, we may abbreviate $\Pr\{f(\mathbf{x})=1\}$, $\Pr\{f(\mathbf{x})=-1\}$ to simply $\Pr\{f\}$, $\Pr\{\neg f\}$ for any binary output functions $f:\mathcal{X}\rightarrow\{-1,+1\}$ when it is necessary for the rest of the paper.
To state the hardness results, we denote $\mathbb{S}^{d-1}:=\{\mathbf{x}\in\mathbb{R}^{d}\ |\ \lVert{\mathbf{x}}\rVert_{2}=1\}$, $\mathbb{Z}_{q}:=\{0,1,\ldots,q-1\}$, $\mathbb{R}_{q}:=[0,q)$, and $\mod_{q}:\mathbb{R}^{d}\rightarrow\mathbb{R}_{q}$ for $q\in\mathbb{N}$. Then we formally define the problem of “learning with errors" (LWE) [25], following [17]:
Definition 2.5 (Learning With Errors).
[…]
  • Alternative hypothesis: $(\mathbf{x},y)$ is generated as $y=\mod_{q}(\mathbf{s}^{\top}\mathbf{x}+z)$, $\mathbf{x}\in\mathcal{D}_{sp},\mathbf{s}\in\mathcal{D}_{st},z\in\mathcal{D}_{ns}$.
  • Null hypothesis: While $\mathbf{x}\in\mathcal{D}_{sp}$, $\mathbf{x}$ and $y$ are independent, where $y$ is sampled uniformly at random on the support of its marginal distribution in alternative hypothesis.
This problem is widely believed to be computationally hard, formalized as follows.
Assumption 2.6 (Sub-exponential LWE Assumption).
[…]
3 FAIRNESS AUDITING AND AGNOSTIC LEARNING
In this section, we describe our reduction from auditing to agnostic learning, which is also our main contribution. In addition, we give a lower bound for fairness auditing under Gaussian distributions.
3.1 Reduction from Auditing to Halfspace Learning
We are considering the auditing problem w.r.t. SPSF as in definition 2.2, which naturally rules out the statistically small subgroups. Indeed, if the probability of accessing the data of certain sub-population is exponentially small, it is computationally hard to even estimate their deviation. Therefore, it makes sense to just consider the collection of subgroups $\mathcal{G}$ that are statistically large enough, e.g., $\Pr\{\mathbf{x}\in g\}=\Theta(1)$ for $\mathbf{x}\in\mathbb{R}^{d}$.
Based on the observation, the following optimization program, $\mathcal{P}_{a,b}^{\mathcal{D}}(\mathcal{H}^{d})$, can capture the most unfair subgroup which is also statistically significant enough. That is
$\max_{g\in\mathcal{G}}\quad$ $\Pr_{\mathbf{x}\in\mathcal{D}}\{\mathbf{x}\in g\}\left|d_{\mathcal{D}}(c,g)\right|$
$s.t.\quad$ $a\leq\Pr_{\mathbf{x}\in\mathcal{D}}\{\mathbf{x}\in g\}\leq b$ (4)
for some constants $10<a\leq b<$ 1.
Furthermore, if we only consider the subgroups represented by halfspaces, there exists a simple reduction from $\mathcal{P}_{a,b}^{\mathcal{D}}(\mathcal{H}^{d})$ to agnostic learning that, in particular, preserves the properties of the data distribution. We show our reduction as the following theorem.
Theorem 3.1 (Main Reduction).
Given any binary classifier $c:\mathbb{R}^{d}\rightarrow\{-1,+1\}$, and a data distribution $\mathcal{D}$ over $\mathbb{R}^{d}$ whose 1-dimensional marginals have continuous cumulative distribution functions, if there exists an efficient algorithm for learning $\mathcal{H}_{\mu}^{\mathcal{D}}$ in the agnostic model on distribution $\mathcal{D}$, then there is an efficient auditing algorithm for $c$ on subgroups represented by $\mathcal{H}^{d}$ over distribution $\mathcal{D}$.
We delay the proof of the above theorem to the end of this section, and show two fundamental hurdles we need to overcome in order to prove theorem 3.1.
Remark 3.2.
While learning from a representation class like $\mathcal{H}_{\mu}^{\mathcal{D}}$ may seems to be hard at a first glance, there are actually examples [23] of learning $\mathcal{H}_{\mu}^{\mathcal{D}}$ in agnostic setting under Gaussian data. Actually, under Gaussian distribution, it is even simpler to consider iteratively learning $\mathcal{H}_{\mu}^{\mathcal{D}}$ with fixed $\mu$, because fixing the threshold allows only optimizing the direction of normal vector due to fact 2.1 (rotational invariance), which resembles to the task of learning homogeneous halfspaces.
Instead of starting from the optimization problem (4), it turns out that solving a sequence of simpler optimization problems suffices to certify the $\gamma$-unfairness as stated in Definition 2.2. We prove the equivalence as the following proposition.
Proposition 3.3.
Consider any binary classifier $c:\mathbb{R}^{d}\rightarrow\{-1,+1\}$, any data distribution $\mathcal{D}$ over $\mathbb{R}^{d}$ whose 1-dimensional marginals have continuous cumulative distribution functions, and any $10<a\leq b<$ 1. For each pair of non-negative integers $k<n$, let $\mathcal{P}_{a,b}^{\mathcal{D}}(k,n)$ denote the optimization program
$\max_{h\in\mathcal{H}^{d}}\quad$ $\Pr_{\mathbf{x}\in\mathcal{D}}\{h(\mathbf{x})=1\}\left|d_{\mathcal{D}}(c,h)\right|$
$s.t.\quad$ $\Pr_{\mathbf{x}\in\mathcal{D}}\{h(\mathbf{x})=1\}=a+k(b-a)/n.$
Let $h^{*}$ be a global optimizer of $\mathcal{P}_{a,b}^{\mathcal{D}}(\mathcal{H}^{d})$, as defined in (4), and let $\gamma^{*}=\Pr\{h^{*}\}\left|d_{\mathcal{D}}(c,h^{*})\right|$. For each $k=0,\ldots,n$, let $h_{k}^{*}$ be a global optimizer of $\mathcal{P}_{a,b}^{\mathcal{D}}(k,n)$. Then
$\max_{k}\Pr\{h_{k}^{*}\}\left|d_{\mathcal{D}}(c,h_{k}^{*})\right|\geq\gamma^{*}-2(b-a)/n.$
Proof.
For conciseness of the proof, we define
$\alpha(k):=a+k(b-a)/n$
Since $a\leq\Pr\{h^{*}(\mathbf{x})=1\}\leq b$ by definition, there must exists a $k\in\{0,\ldots,n-1\}$ such that
$\alpha(k)<\Pr\{h^{*}(\mathbf{x})=1\}<\alpha(k+1)$
Then, since we assumed that $\mathcal{D}$ has a continuous CDF w.r.t. the normal of $h^{*}$, we can construct another halfspace $h^{\prime}$ by either increasing or decreasing the threshold of $h^{*}$ until $\Pr\{\mathbf{x}\in h^{\prime}\}$ hits either $\alpha(k)$ or $\alpha(k+1)$. We thus obtain
$\Pr\{h^{\prime}(\mathbf{x})\neq h^{*}(\mathbf{x})\}=$ $|\Pr\{h^{*}\}-\Pr\{h^{\prime}\}|$
$\leq$ $\alpha(k+1)-\alpha(k)$
$=$ $(b-a)/n$ (5)
Let $\mathbf{dom}:=\{\mathbf{x}\ |\ h^{\prime}(\mathbf{x})\neq h^{*}(\mathbf{x})\}$. Then, by the triangle inequality and the fact that $\Pr\{c(\mathbf{x})=1\}\leq 1$, we have
$\left|\Pr\{h^{*}(\mathbf{x})=1\}d_{\mathcal{D}}(c,h^{*})\right|-\left|\Pr\{h^{\prime}(\mathbf{x})=1\}d_{\mathcal{D}}(c,h^{\prime})\right|\leq$ $|\Pr\{h^{*}\}-\Pr\{h^{\prime}\}|+|\Pr\{h^{\prime}\cap c\}-\Pr\{h^{*}\cap c\}|$
$\overset{\mathrm{}}{\leq}$ $(b-a)/n+|\Pr\{h^{\prime}\cap c\cap\mathbf{dom}\}-\Pr\{h^{*}\cap c\cap\mathbf{dom}\}|$
$\leq$ $(b-a)/n+|\Pr\{\mathbf{x}\in\mathbf{dom}\}|$
$\leq$ $2(b-a)/n$ (6)
where the second inequality is obtained by expanding $\Pr\{h\cap c\}$ on the event $\mathbf{x}\in\mathbf{dom}$ using the law of total probability and exploiting the fact that $h^{\prime}$ always agrees with $h^{*}$ on the complement of $\mathbf{dom}$, i.e., $\Pr\{h^{\prime}\cap c\cap\mathbf{dom}^{c}\}=\Pr\{h^{*}\cap c\cap\mathbf{dom}^{c}\}$; the third inequality holds because at most one of $h^{*}(\mathbf{x})=1$ and $h^{\prime}(\mathbf{x})=1$ holds for any $\mathbf{x}\in\mathbf{dom}$ by definition; and the last inequality is due to equation (5).
Finally, due to the optimality of $h_{k}^{*}$, we have
$\Pr\{h_{k}^{*}\}\left|d_{\mathcal{D}}(c,h_{k}^{*})\right|\geq$ $\Pr\{h^{\prime}\}\left|d_{\mathcal{D}}(c,h^{\prime})\right|-\gamma^{*}+\gamma^{*}$
$\geq$ $\gamma^{*}-2(b-a)/n$
by inequality (6) with $\Pr\{h^{*}(\mathbf{x})=1\}\left|d_{\mathcal{D}}(c,h^{*})\right|=\gamma^{*}$. ∎
The reason why this proposition is so crucial is that it allows us to solve a simpler optimization problem without compromising the guarantee. Being able to fix $\Pr\{h(\mathbf{x})=1\}$ as a constant will significantly simplify the overall optimization as it reduces the degree of the optimization objective. In fact, it is because we can optimize $\Pr\{h(\mathbf{x})=1\}\left|d_{\mathcal{D}}(c,h)\right|$ over $\mathcal{H}_{\mu}^{\mathcal{D}}$ instead of $\mathcal{H}^{d}$ that we can conduct the reduction from auditing to agnostic learning.
The following lemma shows a direct relationship between the unfairness level and the classification error.
Lemma 3.4.
Given any binary classifier $c:\mathcal{X}\rightarrow\{-1,+1\}$, a data distribution $\mathcal{D}$ over $\mathcal{X}$ and a collection of subgroups $g\in\mathcal{G}$ such that $g:\mathcal{X}\rightarrow\{-1,+1\}$, we have
$2\Pr\{g\}d_{\mathcal{D}}(c,g)=\Pr\{\neg c\}\Pr\{\neg g\}+\Pr\{c\}\Pr\{g\}-\Pr\{c(\mathbf{x})=g(\mathbf{x})\}$
for $\mathbf{x}\sim\mathcal{D}$.
Proof.
By the law of total probability, we have
$\Pr\{c\cap g\}=\Pr\{g\}-(\Pr\{\neg c\}-\Pr\{\neg c\cap\neg g\}).$
which along with definition 2.2 gives
$d_{\mathcal{D}}(c,g)=$ $\Pr\{c\}-\Pr\{c\ |\ g\}$
$=$ $\frac{\Pr\{c\}\Pr\{g\}-\Pr\{c\cap g\}}{\Pr\{g\}}$
$=$ $\frac{\Pr\{\neg c\}\Pr\{\neg g\}-\Pr\{\neg c\cap\neg g\}}{\Pr\{g\}}.$ (7)
Summing up the two different forms of $d_{\mathcal{D}}(c,g)$ results to
$2d_{\mathcal{D}}(c,g)=$ $\frac{\Pr\{\neg c\}\Pr\{\neg g\}-\Pr\{\neg c\cap\neg g\}}{\Pr\{g\}}+\frac{\Pr\{c\}\Pr\{g\}-\Pr\{c\cap g\}}{\Pr\{g\}}$
$=$ $\frac{\Pr\{\neg c\}\Pr\{\neg g\}+\Pr\{c\}\Pr\{g\}-(\Pr\{\neg c\cap\neg g\}+\Pr\{c\cap g\})}{\Pr\{g\}}$ (8)
Notice that, because $c\cap g$ and $\neg c\cap\neg g$ are two disjoint events, we have
$\Pr\{c(\mathbf{x})=g(\mathbf{x})\}=$ $\Pr\{(c\cap g)\cup(\neg c\cap\neg g)\}$
$=$ $\Pr\{c\cap g\}+\Pr\{\neg c\cap\neg g\}$
Plugging it back to equation (8) produces the desired result. ∎
This immediately implies the duality between SPSF auditing and agnostic learning as follow.
Corollary 3.5.
Given any binary classifier $c:\mathbb{R}^{d}\rightarrow\{-1,+1\}$, a data distribution $\mathcal{D}$ and a collection of halfspaces $\mathcal{H}_{\mu}^{\mathcal{D}}$ over $\mathbb{R}^{d}$, we have the following two properties
  1. (1)
    $d_{\mathcal{D}}(c,h^{*})\geq d_{\mathcal{D}}(c,h),\forall h\in\mathcal{H}_{\mu}^{\mathcal{D}}$ if and only if $h^{*}=\operatorname*{argmin}_{h\in\mathcal{H}_{\mu}^{\mathcal{D}}}\Pr_{\mathbf{x}\sim\mathcal{D}}\{c(\mathbf{x})=h(\mathbf{x})\}$
  2. (2)
    $d_{\mathcal{D}}(c,h^{*})\leq d_{\mathcal{D}}(c,h),\forall h\in\mathcal{H}_{\mu}^{\mathcal{D}}$ if and only if $h^{*}=\operatorname*{argmax}_{h\in\mathcal{H}_{\mu}^{\mathcal{D}}}\Pr_{\mathbf{x}\sim\mathcal{D}}\{c(\mathbf{x})=h(\mathbf{x})\}$
Proof.
Because $\Pr\{c\}$ is a constant and $\Pr\{h\}=\mu,\forall h\in\mathcal{H}_{\mu}^{\mathcal{D}}$ by Definition 2.4, $d_{\mathcal{D}}(c,h)$ is simply an affine transformation of $\Pr\{c(\mathbf{x})=h(\mathbf{x})\}$ for a fixed $\mu$ by Lemma 3.4, which implies the desired results. ∎
Proposition 3.3 tells us that solving $\mathcal{P}_{a,b}^{\mathcal{D}}(k,n)$ for $k=0,\ldots,n$ would give us a good enough approximation to the maximum unfairness level, of course, with a large enough $n$. Therefore, we just need to further show that solving each $\mathcal{P}_{a,b}^{\mathcal{D}}(k,n)$ is equivalent to learning $\mathcal{H}_{\mu}^{\mathcal{D}}$ to complete the reduction.
Formally, because $\mathcal{P}_{a,b}^{\mathcal{D}}(k,n)$ can be equivalently written as
$\max_{h\in\mathcal{H}_{\mu}^{\mathcal{D}}}\quad\Pr_{\mathbf{x}\in\mathcal{D}}\{h(\mathbf{x})=1\}\left|d_{\mathcal{D}}(c,h)\right|$ (9)
for some $\mu=a+k(b-a)/n$, it suffices to prove the following theorem.
Lemma 3.6.
Given any binary classifier $c:\mathbb{R}^{d}\rightarrow\{-1,+1\}$, a data distribution $\mathcal{D}$ and a collection of halfspaces $\mathcal{H}_{\mu}^{\mathcal{D}}$ over $\mathbb{R}^{d}$ such that
$\mathrm{opt}_{\min}\leq\Pr_{\mathbf{x}\sim\mathcal{D}}\{c(\mathbf{x})=h(\mathbf{x})\}\leq\mathrm{opt}_{\max}$
for all $h\in\mathcal{H}_{\mu}^{\mathcal{D}}$, if $h_{\mathbf{v}},h_{\mathbf{u}}\in\mathcal{H}_{\mu}^{\mathcal{D}}$ satisfy that $\Pr\{c(\mathbf{x})=h_{\mathbf{v}}(\mathbf{x})\}\leq\mathrm{opt}_{\min}+2\epsilon$ as well as $\Pr\{c(\mathbf{x})=h_{\mathbf{u}}(\mathbf{x})\}\geq\mathrm{opt}_{\max}-2\epsilon$, we have either
$\Pr_{\mathbf{x}\sim\mathcal{D}}\{h_{\mathbf{v}}(\mathbf{x})=1\}\left|d_{\mathcal{D}}(c,h_{\mathbf{v}})\right|\geq\gamma^{*}-\epsilon$ (10)
or
$\Pr_{\mathbf{x}\sim\mathcal{D}}\{h_{\mathbf{u}}(\mathbf{x})=1\}\left|d_{\mathcal{D}}(c,h_{\mathbf{u}})\right|\geq\gamma^{*}-\epsilon$ (11)
where $\gamma^{*}=\max_{h\in\mathcal{H}_{\mu}^{\mathcal{D}}}\Pr_{\mathbf{x}\sim\mathcal{D}}\{h(\mathbf{x})=1\}\left|d_{\mathcal{D}}(c,h)\right|$.
Proof.
By the proof of Lemma 3.4, we have
$2\Pr\{h\}\left|d_{\mathcal{D}}(c,h)\right|=|\underbrace{\Pr\{\neg c\}\Pr\{\neg h\}-\Pr\{\neg c\cap\neg h\}}_{I_{1}}+\underbrace{\Pr\{c\}\Pr\{h\}-\Pr\{c\cap h\}}_{I_{2}}|$
Let $h^{*}\in\mathcal{H}_{\mu}^{\mathcal{D}}$ be such that $\Pr\{h^{*}\}\left|d_{\mathcal{D}}(c,h^{*})\right|=\gamma^{*}$. Then for $I_{2}$, we have
$I_{2}=$ $(\Pr\{c\}-\Pr\{c\ |\ h^{*}\}+\Pr\{c\ |\ h^{*}\})\Pr\{h\}-\Pr\{c\cap h\}$
$=$ $\Pr\{h^{*}\}d_{\mathcal{D}}(c,h^{*})+\Pr\{c\cap h^{*}\}-\Pr\{c\cap h\}$
where the last equation is because $h^{*}\in\mathcal{H}_{\mu}^{\mathcal{D}}$, then $\Pr\{h\}=\Pr\{h^{*}\}=\mu$ by Definition 2.4.
Similarly, for $I_{1}$, we can write
$I_{1}=$ $\Pr\{\neg h^{*}\}(\Pr\{\neg c\}-\Pr\{\neg c\ |\ \neg h^{*}\})+\Pr\{\neg c\cap\neg h^{*}\}-\Pr\{\neg c\cap\neg h\}$
$=$ $\Pr\{h^{*}\}d_{\mathcal{D}}(c,h^{*})+\Pr\{\neg c\cap\neg h^{*}\}-\Pr\{\neg c\cap\neg h\}$
where the last equation follows because we have shown in the proof of Lemma 3.4 that $d_{\mathcal{D}}(c,h^{*})=\Pr\{\neg h^{*}\}(\Pr\{\neg c\}-\Pr\{\neg c\ |\ \neg h^{*}\})/\Pr\{h^{*}\}$.
Combining $I_{1}$ and $I_{2}$ will result to
$\Pr\{h\}\left|d_{\mathcal{D}}(c,h)\right|=$ $|\Pr\{h^{*}\}d_{\mathcal{D}}(c,h^{*})+(\Pr\{c(\mathbf{x})=h^{*}(\mathbf{x})\}-\Pr\{c(\mathbf{x})=h(\mathbf{x})\})/2|$
$\geq$ $\gamma^{*}-\left|\Pr\{c(\mathbf{x})=h^{*}(\mathbf{x})\}-\Pr\{c(\mathbf{x})=h(\mathbf{x})\}\right|/2$
by triangle inequality. Further, since $h^{*}$ maximizes $\left|d_{\mathcal{D}}(c,h)\right|$, it either maximizes or minimizes $d_{\mathcal{D}}(c,h)$. Then, by Corollary 3.5, we know
$\Pr_{\mathbf{x}\sim\mathcal{D}}\{c(\mathbf{x})=h^{*}(\mathbf{x})\}\in\{\mathrm{opt}_{\min},\mathrm{opt}_{\max}\}$
which implies either
$\left|\Pr\{c(\mathbf{x})=h^{*}(\mathbf{x})\}-\Pr\{c(\mathbf{x})=h_{\mathbf{v}}(\mathbf{x})\}\right|/2\leq\epsilon$
or
$\left|\Pr\{c(\mathbf{x})=h^{*}(\mathbf{x})\}-\Pr\{c(\mathbf{x})=h_{\mathbf{u}}(\mathbf{x})\}\right|/2\leq\epsilon$
Therefore, the proof is completed. ∎
Remark 3.7.
We emphasize that it is necessary for us to consider the guarantee of agnostic learning in a additive form rather than multiplicative form. Although Corollary 3.5 shows that the classification error, $\Pr\{c(\mathbf{x})\neq h(\mathbf{x})\}$, and the unfairness level, $\Pr\{h\}\left|d_{\mathcal{D}}(c,h)\right|$, are dual to each other over $\mathcal{H}_{\mu}^{\mathcal{D}}$, the affine relationship between them prohibits obtaining a guarantee on the unfairness from a multiplicative error. This also explains why the guarantee provided by [23] does not fit in our analysis.
Now we are ready to prove Theorem 3.1.
Proof of Theorem 3.1.
To solve the auditing problem, we just need to solve the sequence of optimization problems, $\{\mathcal{P}_{a,b}^{\mathcal{D}}(k,n)\ |\ k=0,\ldots,n\}$ as described in Proposition 3.3. We can solve each $\mathcal{P}_{a,b}^{\mathcal{D}}(k,n)$ with an additive error $\epsilon$ by calling the given oracle of learning halfspaces with the same strategy specified in Lemma 3.6. Eventually, we solve all of these optimization problems with an $2(b-a)/n+\epsilon$ additive error and a running time of $O(n)$ factor overhead compared with that of the oracle. ∎
3.2 The Hardness of Auditing Under Gaussian Data
We show that the problem of auditing halfspaces subgroups under Gaussian distribution is computationally hard in two forms: the multiplicative form and additive form. The hardness results of both cases are obtained through reduction from auditing halfspace subgroups to the problem of continuous Learning With Error (cLWE) under Gaussian distribution, which is known to be as hard as LWE.
Proposition 3.8 ([26, 17] Hardness of cLWE).
[…]
3.2.1 Approximation in Multiplicative Form
Suppose an auditing algorithm is guaranteed to return us a $\gamma^{\prime}$-unfair certificate (a halfspace) give a $\gamma$-unfair classifier $c$, where $\gamma^{\prime}\leq\gamma\leq 1$, we can show that $\gamma^{\prime}$ can never be close to $\gamma$. We can show that $\gamma^{\prime}=\Omega(\gamma/\mathrm{poly}(d))$ only if cLWE can be solved in $O(\mathrm{poly}(d))$ time.
Theorem 3.9.
Under Assumption 2.6, for any $d\in\mathbb{N}$, any constants $\alpha\in(0,1),\beta\in\mathbb{R}_{+}$, and any $\log^{\beta}d\leq k\leq cd$ where $c$ is a sufficiently small constant, there is not algorithm that runs in time $d^{O(k^{\alpha})}$ and distinguishes between the following two cases of a joint distribution $\mathcal{D}$ of $(\mathbf{x},c(\mathbf{x}))$ supported on $\mathbb{R}^{d}\times\{-1,+1\}$ with marginal $\mathcal{D}_{\mathbf{x}}=\mathcal{N}(0,\mathrm{I})$, with $d^{-O(k^{\alpha})}$ advantage:
  • (i)
    Alternative Hypothesis: There exists non-negligible unfair halfspace subgroups, i.e., $\exists h\in\mathcal{H}^{d},\Pr_{\mathcal{D}}\{h(\mathbf{x})=1\}|d_{\mathcal{D}}(c,h)|=\Omega(1/\sqrt{k\log d})$.
  • (ii)
    Null Hypothesis: All halfspace subgroups are absolutely fair, i.e., $\Pr_{\mathcal{D}}\{h(\mathbf{x})=1\}|d_{\mathcal{D}}(c,h)|=0,\forall h\in\mathcal{H}^{d}$.
Proof.
[…]
$c(\mathbf{x})=\begin{cases}+1,\quad\text{if }y\leq T/2\\ -1,\quad\text{otherwise}\end{cases}$
We argue that $\mathcal{D}$ satisfies the desired requirement stated above.
For the alternative hypothesis case, let $\mathcal{D}^{\prime}$ be from the alternative hypothesis case of the LWE. Let $\mathbf{s}$ be the secret vector in the LWE problem. We consider the following two halfspaces:
$h_{1}(\mathbf{x})=$ $\operatorname*{\mathrm{sgn}}(\mathbf{s}^{\top}\mathbf{x}-T/6)$
$h_{2}(\mathbf{x})=$ $\operatorname*{\mathrm{sgn}}(-\mathbf{s}^{\top}\mathbf{x}+T/3)$
If we can show $\left|\Pr_{\mathbf{x}\sim\mathcal{D}_{\mathbf{x}}}\{h_{1}(\mathbf{x})=1\}d_{\mathcal{D}}(c,h_{1})+\Pr_{\mathbf{x}\sim\mathcal{D}_{\mathbf{x}}}\{h_{2}(\mathbf{x})=1\}d_{\mathcal{D}}(c,h_{2})\right|=\Omega(T)$, then either $h=h_{1}$ or $h=h_{2}$ satisfies $\Pr_{\mathbf{x}\sim\mathcal{D}_{\mathbf{x}}}\{h(\mathbf{x})=1\}\left|d_{\mathcal{D}}(c,h)\right|=\Omega(T)$, which implies the desired property of the alternative hypothesis we would like to prove. By lemma 3.4, we have
$2\Pr_{\mathbf{x}\sim\mathcal{D}_{\mathbf{x}}}\{h_{1}(\mathbf{x})=1\}d_{\mathcal{D}}(c,h_{1})+2\Pr_{\mathbf{x}\sim\mathcal{D}_{\mathbf{x}}}\{h_{2}(\mathbf{x})=1\}d_{\mathcal{D}}(c,h_{2})$
$=$ $\underbrace{\Pr\{\neg c\}(\Pr\{\neg h_{1}\}+\Pr\{\neg h_{2}\})+\Pr\{c\}(\Pr\{h_{1}\}+\Pr\{h_{2}\})}_{I_{1}}-(\underbrace{\Pr\{c(\mathbf{x})\neq h_{1}(\mathbf{x})\}+\Pr\{c(\mathbf{x})\neq h_{2}(\mathbf{x})\}}_{I_{2}})$
To bound $I_{1},I_{2}$, we first examine the subset of domain where $h_{1}$ and $h_{2}$ agree, namely
$B:=$$\{\mathbf{x}\in\mathbb{R}^{d}\ |\ h_{1}(\mathbf{x})=h_{2}(\mathbf{x})\}$
$=$ $\{\mathbf{x}\in\mathbb{R}^{d}\ |\ h_{1}(\mathbf{x})=1\ \cap\ h_{2}(\mathbf{x})=1\}$
$=$ $\{\mathbf{x}\in\mathbb{R}^{d}\ |\ \mathbf{s}^{\top}\mathbf{x}\in[T/6,T/3]\}$
Then, for $I_{1}$, by the law of total probability, we have
$I_{1}=$ $\Pr\{c(\mathbf{x})=-1\}(\Pr\{h_{1}(\mathbf{x})=-1\}+\Pr\{h_{2}(\mathbf{x})=-1\}+\Pr\{\mathbf{x}\in B\}-\Pr\{\mathbf{x}\in B\})$
$+\Pr\{c(\mathbf{x})=1\}(\Pr\{h_{1}(\mathbf{x})=1\}+\Pr\{h_{2}(\mathbf{x})=1\ \cap\ \mathbf{x}\notin B\}+\Pr\{h_{2}(\mathbf{x})=1\ \cap\ \mathbf{x}\in B\})$
=$\overset{\mathrm{(i)}}{=}$ $\Pr\{c(\mathbf{x})=-1\}(1-\Pr\{\mathbf{x}\in B\})+\Pr\{c(\mathbf{x})=1\}(1+\Pr\{\mathbf{x}\in B\})$
$=$ $1+\Pr\{\mathbf{x}\in B\}(\Pr\{c(\mathbf{x})=1\}-\Pr\{c(\mathbf{x})=-1\})$
where (i) is because $\{\mathbf{x}\in\mathbb{R}^{d}\ |\ h_{1}(\mathbf{x})=-1\},\{\mathbf{x}\in\mathbb{R}^{d}\ |\ h_{2}(\mathbf{x})=-1\},\{\mathbf{x}\in B\}$ are pairwise disjoint and their union equals to $\mathbb{R}^{d}$, $\{\mathbf{x}\in\mathbb{R}^{d}\ |\ h_{1}(\mathbf{x})=1\},\{\mathbf{x}\in\mathbb{R}^{d}\ |\ h_{2}(\mathbf{x})=1\ \cap\ \mathbf{x}\notin B\}$ are disjoint and their union equals to $\mathbb{R}^{d}$, also, $\{\mathbf{x}\in B\}\subset\{\mathbf{x}\in\mathbb{R}^{d}\ |\ h_{2}(\mathbf{x})=1\}$ by definition.
For $I_{2}$, because for any $\mathbf{x}\in B$, $h_{1}(\mathbf{x})=h_{2}(\mathbf{x})=1$ by construction, and by the law of total probability, we have
$I_{2}=$ $\Pr\{c(\mathbf{x})\neq h_{1}(\mathbf{x})\ \cap\ \mathbf{x}\notin B\}+\Pr\{c(\mathbf{x})\neq h_{2}(\mathbf{x})\ \cap\ \mathbf{x}\notin B\}+2\Pr\{c(\mathbf{x})=-1\ \cap\ \mathbf{x}\in B\}$
$=$ $\Pr\{\mathbf{x}\notin B\}+2\Pr\{c(\mathbf{x})=-1\ \cap\ \mathbf{x}\in B\}$
$=$ $1+\Pr\{c(\mathbf{x})=-1\ \cap\ \mathbf{x}\in B\}-\Pr\{c(\mathbf{x})=1\ \cap\ \mathbf{x}\in B\}$
$=$ $1-\Pr\{\mathbf{x}\in B\}(1-2\Pr\{c(\mathbf{x})=-1\ |\ \mathbf{x}\in B\}$
By the definition of $c$ as well as the Alternative case distribution of the LWE problem, $\{\mathbf{x}\in\mathbb{R}^{d}\ |\ c(\mathbf{x})=1\}$ is equivalent to $\{\mathbf{x}\in\mathbb{R}^{d}\ |\ \mod_{T}(\mathbf{s}^{\top}\mathbf{x}+z)\leq T/2\}$ for some $z\sim\mathcal{N}(0,\sigma^{2})$. Further, we have
$\{\mathbf{x}\in\mathbb{R}^{d}\ |\ \mod_{T}(\mathbf{s}^{\top}\mathbf{x}+z)\leq T/2\}\equiv\ \cup\ _{k\in\mathbb{Z}}\{\mathbf{s}^{\top}\mathbf{x}+z\in(kT,kT+T/2]\}$
Notice that $\mathbf{s}^{\top}\mathbf{x}+z$ is a one dimensional Gaussian random variable, which, by symmetry of Gaussian distribution, implies $\Pr\{c(\mathbf{x})=1\}=\Pr\{\ \cup\ _{k\in\mathbb{Z}}\{\mathbf{s}^{\top}\mathbf{x}+z\in(kT,kT+T/2]\}\}=1/2$. Therefore, combining $I_{1}$ and $I_{2}$ gives
$I_{1}-I_{2}=$ $\Pr\{\mathbf{x}\in B\}(1-2\Pr\{c(\mathbf{x})=-1\ |\ \mathbf{x}\in B\})$
$=$ $\Omega(T)(1-2\Pr\{c(\mathbf{x})=-1\ |\ \mathbf{x}\in B\})$ (12)
where the last equation is because $\mathbf{s}^{\top}\mathbf{x}\sim\mathcal{N}(0,1)$, hence, $\Pr\{\mathbf{x}\in B\}=\Pr\{\mathbf{s}^{\top}\mathbf{x}\in[T/6,T/3]\}=\Omega(T)$. Thus, it suffices to show $1-2\Pr\{c(\mathbf{x})=-1\ |\ \mathbf{x}\in B\}=\Omega(1)$.
Again, by the definition of the Alternative case distribution from the LWE problem, we know that $c(\mathbf{x})=-1$ only if $\mod_{T}(\mathbf{s}^{\top}\mathbf{x}+z)>T/2$, which in turn happens only if $\mathbf{s}^{\top}\mathbf{x}+z>T/2$ or $\mathbf{s}^{\top}\mathbf{x}+z<0$. For $\mathbf{x}\in B$, we have $\mathbf{s}^{\top}\mathbf{x}\in[T/6,T/3]$, therefore $c(\mathbf{x})=-1$ only if $\left|z\right|\geq T/6$. Notice that $z\sim\mathcal{N}(0,\sigma^{2})$ and Proposition 3.8 states that the LWE problem is hard for any fixed constant $\kappa\in\mathbb{N}$ and $\sigma\geq k^{-\kappa}$. Given the constant $\beta\in\mathbb{R}_{+}$ in this theorem, we can take $\kappa=\left\lceil 1/2\beta+1/2+1\right\rceil$, which is a fixed constant. Then, by Proposition 3.8, the LWE problem is hard for $\sigma=k^{-\kappa}\leq 1(k^{3/2}\sqrt{\log d})=o(T)$. Therefore, by Gaussian tail bound, we have
$\Pr_{\mathbf{x}\sim\mathcal{D}_{\mathbf{x}}}\{c(\mathbf{x})=-1\ |\ \mathbf{x}\in B\}\leq\Pr_{z\sim\mathcal{N}(0,\sigma^{2})}\{\left|z\right|\geq T/6\}=o(1)$
Plugging the above back to equation (12), we can conclude that
$\Pr_{\mathbf{x}\sim\mathcal{D}_{\mathbf{x}}}\{h_{1}(\mathbf{x})=1\}d_{\mathcal{D}}(c,h_{1})+\Pr_{\mathbf{x}\sim\mathcal{D}_{\mathbf{x}}}\{h_{2}(\mathbf{x})=1\}d_{\mathcal{D}}(c,h_{2})=\Omega(T)$
Thus, either $h=h_{1}$ or $h=h_{2}$ must satisfies $\Pr_{\mathbf{x}\sim\mathcal{D}_{\mathbf{x}}}\{h(\mathbf{x})=1\}\left|d_{\mathcal{D}}(c,h)\right|=\Omega(T)$, which completes the proof for the alternative hypothesis case.
For the null hypothesis, we can immediately see that $\Pr_{\mathbf{x}\in\mathcal{N}}\{h\}d_{\mathcal{N}}(c,h)=0,\forall h\in\mathcal{H}^{d}$ because $c(\mathbf{x})$ is independent from each $h\in\mathcal{H}^{d}$.
[…]
It immediately indicates the following corollary.
Corollary 3.10.
Given Assumption 2.6, for any constants $\alpha\in(0,1),\log^{\beta}d\leq k\leq cd$ where $c$ is a sufficiently small constant, there is no polynomial-time $1/\mathrm{poly}(d)$-approximation algorithms for auditing halfspace subgroups under Gaussian marginal.
Proof.
Suppose there exists an auditing algorithm that guarantees to return a $\delta\gamma$-unfair certificate given a $\gamma$-unfair collection of halfspace subgroup and access to data with Gaussian marginal, where $\delta\in(0,1)$.
For the alternative hypothesis case as described in Theorem 3.9, given a $1/\sqrt{k\log d}$-unfair collection of halfspace subgroups, we run such an algorithm to obtain a $\delta/\sqrt{k\log d}$-unfair certificate, i.e., a halfspace $h$ such that $\Pr_{\mathbf{x}\sim\mathcal{N}}\{h(\mathbf{x})=1\}|d_{\mathcal{N}}(c,h)|=\Omega(\delta/\sqrt{k\log d})$. Notice that by Hoeffding Bound, with $d^{O(k^{\alpha})}$ examples from the distribution constructed in the alternative hypothesis case, we can verify that the empirical estimation of $\Pr_{\mathbf{x}\sim\mathcal{N}}\{h(\mathbf{x})=1\}|d_{\mathcal{N}}(c,h)|$ is $d^{-O(k^{\alpha})}$-close to $\Omega(\delta/\sqrt{k\log d})$ with high probability.
For the null hypothesis case, with the same argument, we can verify there is no $d^{-O(k^{\alpha})}$-unfair subgroup with high probability given $d^{O(k^{\alpha})}$ examples from the distribution in the null hypothesis case.
Therefore, the algorithm can distinguish the alternative hypothesis case from the null hypothesis case in Theorem 3.9 unless $\delta=d^{-O(k^{\alpha})}$, which implies the desired result. ∎
3.2.2 Approximation in Additive Form
According to Theorem 3.1, how well we can approximate the most unfair subgroup is controlled by the parameter $\epsilon$. Unfortunately, it turns out, even under a Gaussian distribution, $\epsilon$ cannot be as small as $1/\mathrm{poly}(d)$.
A recent result [17] showed that learning halfspaces with an additive error $O(1/\sqrt{\log d})$ is as least as hard as the problem of continuous Learning With Error (cLWE) under Gaussian distribution.
In particular, they showed an efficient way to generate $(\mathbf{x},y)\in\mathbb{R}^{d}\times\{-1,+1\}$ from the data, $(\mathbf{x},y^{\prime})\in\mathbb{R}^{d}\times\mathbb{R}_{T}$, sampled from the cLWE structured Gaussian distribution such that the properties of both alternative and null hypothesis are preserved. In particular, under the distribution from the alternative hypotheses of cLWE, they claim that at least one of the two halfspaces, $\mathbf{s}^{\top}\mathbf{x}\geq T/6$ and $\mathbf{s}^{\top}\mathbf{x}\leq T/3$, must have classification error less than $1/2-\Omega(T)$. Hence, if one can learn halfspaces, they can also distinguish the two cases in cLWE.
We can reduce cLWE to auditing by generating data in the same way as [17]. For the null hypothesis, we have that $\Pr_{\mathbf{x}\in\mathcal{N}}\{h\}d_{\mathcal{N}}(c,h)=0,\forall h\in\mathcal{H}^{d}$ because $c(\mathbf{x})$ is independent from each $h\in\mathcal{H}^{d}$. For the alternative hypothesis, the same two halfspaces, $\mathbf{s}^{\top}\mathbf{x}\geq T/6$ and $\mathbf{s}^{\top}\mathbf{x}\leq T/3$, will result to a unfairness level distinguishable from 00 based on Lemma 3.6. Therefore, being able to audit $h_{1}$ and $-h_{2}$ such that $h_{1}\in\mathcal{H}_{T/6}^{\mathcal{N}}$ and $h_{2}\in\mathcal{H}_{-T/3}^{\mathcal{N}}$ to an $\epsilon$ error will allow us to solve cLWE problem. This observation yields the following corollary.
Corollary 3.11.
Given Assumption 2.6, for any constants $\alpha\in(0,2),\beta>1/2$ and any $C/\sqrt{d\log d}\leq\epsilon\leq 1/\log^{\beta}d$ where $C$ is a sufficiently large constant, there is no algorithm running in time $d^{O(\epsilon\sqrt{\log d})^{\alpha}}$ that can recover a subgroup represented by a halfspace on $\mathbb{R}^{d}$ with Gaussian marginals with unfairness within an additive $\epsilon$ of maximum.
For details about the reduction, we refer to the proof of [17].
To the best of our knowledge, there does not exist any PTAS approach for properly learning general halfspaces in agnostic model with guarantees of additive error close to $O(1/\sqrt{\log d})$. However, in the next section, we will show that if we restrict out attention to just homogeneous halfspaces under a standard normal distribution, it is possible to achieve additive error of $O(1/\log^{1/C}d)$ for some constant $C>2$.
4 AUDITING VIA AGNOSTIC LEARNING UNDER GAUSSIAN DISTRIBUTION
In this section, we present our algorithmic results. Our approach is based on Theorem 3.1: auditing over subgroups determined by halfspaces can be accomplished by solving a sequence of simpler tasks of learning halfspaces. As a result, we are able to take advantage of existing agnostic learning methods to solve the auditing problem.
Assuming there exists an efficient oracle for agnostic learning, algorithm 1 will eventually return a halfspace $h^{\prime}$ as a certificate of the subgroup that has the highest unfairness level. We show the correctness, time and sample complexity of algorithm 1 in Theorem 4.1.
Input : $n,a,b,\epsilon,\delta,\mathcal{D}$, classifier $c$, oracle $\mathcal{O}$
Result: $\mu^{\prime},h^{\prime}$
1 $\hat{\mathcal{X}}\leftarrow$ draw $N(d,\epsilon,\delta)$ i.i.d. samples from $\mathcal{D}$;
2 $\hat{\mathcal{D}^{+}}\leftarrow\{\hat{\mathcal{X}},c(\hat{\mathcal{X}})\}$;
3 $\hat{\mathcal{D}^{-}}\leftarrow\{\hat{\mathcal{X}},-c(\hat{\mathcal{X}})\}$;
4 $\mu\leftarrow a$;
5 $(\mu^{\prime},h^{\prime})\leftarrow(1,c)$;
6 while $\mu\leq b$ do
7       $h_{\mu}^{+}\leftarrow\mathcal{O}(\epsilon,\delta/2n,\mu,\hat{\mathcal{D}^{+}})$;
8       $h_{\mu}^{-}\leftarrow\mathcal{O}(\epsilon,\delta/2n,\mu,\hat{\mathcal{D}^{-}})$;
9       if $\left|d_{\mathcal{D}}(c,h_{\mu}^{+})\right|<\left|d_{\mathcal{D}}(c,h_{\mu}^{-})\right|$ then
10             $h_{\mu}^{+}\leftarrow h_{\mu}^{-}$;
11
12       end if
13      if $\mu^{\prime}\left|d_{\mathcal{D}}(c,h^{\prime})\right|\leq\mu\left|d_{\mathcal{D}}(c,h_{\mu}^{+})\right|$ then
14             $(\mu^{\prime},h^{\prime})\leftarrow(\mu,h_{\mu}^{+})$;
15
16       end if
17      $\mu\leftarrow\mu+(b-a)/n$;
18
19 end while
Algorithm 1 Fairness Auditing
Theorem 4.1 (Auditing Framework).
Given any binary classifier $c:\mathbb{R}^{d}\rightarrow\{-1,+1\}$, a data distribution $\mathcal{D}$ whose 1-dimensional marginals have continuous cumulative distribution functions, and collections of halfspaces $\{\mathcal{H}_{\mu}^{\mathcal{D}}\ |\ \mu>0\}$ over $\mathbb{R}^{d}$, if there exists an oracle $\mathcal{O}$ that takes $\epsilon,\delta\,\mu\in(0,1)$ and $N(d,\epsilon,\delta)$ labeled i.i.d. samples from $\mathcal{D}$ in the form of $(\mathbf{x},c(\mathbf{x}))$, runs in time $T(d,\epsilon,\delta)$, and returns a halfspace $h_{\mu}$ such that, with at least $1-\delta$ probability
$\Pr_{\mathbf{x}\sim\mathcal{D}}\{h_{\mu}(\mathbf{x})\neq c(\mathbf{x})\}\leq\min_{h\in\mathcal{H}_{\mu}^{\mathcal{D}}}\Pr_{\mathbf{x}\sim\mathcal{D}}\{h(\mathbf{x})\neq c(\mathbf{x})\}+\epsilon$
then there exists an algorithm that takes $n\in\mathbb{Z}^{+}$, $10<a\leq b<$ 1, $\epsilon,\delta\in(0,1)$ and $O(N(d,\epsilon,\delta/n))$ labeled i.i.d. samples from $\mathcal{D}$, runs in time $O(nT(d,\epsilon,\delta/n))$, returns a halfspace $h^{\prime}$ as a certificate such that $a\leq\Pr_{\mathbf{x}\sim\mathcal{D}}\{h^{\prime}\}\leq b$ as well as
$\Pr_{\mathbf{x}\sim\mathcal{D}}\{h^{\prime}\}\left|d_{\mathcal{D}}(c,h^{\prime})\right|\geq\max_{h\in\mathcal{H}^{d}}\Pr_{\mathbf{x}\sim\mathcal{D}}\{h\}\left|d_{\mathcal{D}}(c,h)\right|-O(\epsilon)$
with at least $1-\delta$ probability.
Proof.
Let’s notice that, although each iteration of the loop in algorithm 1 solves $\min_{h\in\mathcal{H}_{\mu}^{\mathcal{D}}}\Pr\{c(\mathbf{x})\neq h(\mathbf{x})\}$ and $\max_{h\in\mathcal{H}_{\mu}^{\mathcal{D}}}\Pr\{c(\mathbf{x})\neq h(\mathbf{x})\}$, it is essentially equivalent to solving $\max_{h\in\mathcal{H}_{\mu}^{\mathcal{D}}}\left|d_{\mathcal{D}}(c,h)\right|$ according to lemma 3.6. As the oracle returns a halfspace with additive error smaller than $\epsilon$ with probability at least $1-\delta$, we have that
$\max(\left|d_{\mathcal{D}}(c,h_{\mu}^{+})\right|,\left|d_{\mathcal{D}}(c,h_{\mu}^{-})\right|)\geq\max_{h\in\mathcal{H}_{\mu}^{\mathcal{D}}}\left|d_{\mathcal{D}}(c,h_{\mu}^{+})\right|-\frac{\epsilon}{\mu}$
with probability at least $1-\delta/n$ because of Lemma 3.6 as well as union bound.
Across all iterations, the algorithm maximizes $\mu\left|d_{\mathcal{D}}(c,h_{\mu}^{+})\right|$ over $\mathcal{H}_{\mu}^{\mathcal{D}}$ for $\mu$ increase from $a$ to $b$ with step size $(b-a)/n$. With a union bound over all $n$ iterations, we can have the same additive error $\epsilon$ in every iterations, with probability at least $1-\delta$. As a result, the algorithm equivalently solves
$\max_{h\in\mathcal{H}^{d}}\quad$ $\Pr_{\mathbf{x}\in\mathcal{D}}\{h(\mathbf{x})=1\}\left|d_{\mathcal{D}}(c,h)\right|$
$s.t.\quad$ $a\leq\Pr_{\mathbf{x}\in\mathcal{D}}\{h(\mathbf{x})=1\}\leq b$
with probability at least $1-\delta$ for an additive error at most $2(b-a)/n+\epsilon$ according to Proposition 3.3, which completes the proof. ∎
While our framework heavily relies on the methods of agnostic learning with small additive error, unfortunately, there are no known methods for learning general halfspaces that can achieve additive error better than a constant, even under distributions as nice as standard normal ones.
However, if we restrict our auditing subjects to the class of homogeneous halfspaces, [18] proposed an agnostic learning PTAS for homogeneous halfspaces under Gaussian data. That is, we only audit for subgroups with probability mass $1/2$.
Lemma 4.2 (Learning Homogeneous Halfspaces [18]).
Let $\mathcal{D}$ be a distribution on labeled examples $(\mathbf{x},y)\in\mathbb{R}^{d}\times\{-1,+1\}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(0,\mathrm{I})$. There exists an algorithm that, given $\epsilon,\delta>0$, and $N=d^{O(1/\epsilon^{4})}\mathrm{poly}(1/\epsilon)\log(1/\delta)$ i.i.d. samples from $\mathcal{D}$, the algorithm runs in time $\mathrm{poly}(N)+(1/\epsilon)^{O(1/\epsilon^{6})}\log(1/\delta)$, and computes a homogeneous halfspace $h_{\mathbf{v}}$ such that, with probability at least $1-\delta$, it holds $\Pr_{\mathcal{D}}\{y\neq h_{\mathbf{v}}(\mathbf{x})\}\leq\mathrm{opt}+\epsilon$.
We show our algorithmic guarantee of a PTAS in the following corollary.
Corollary 4.3 (Auditing Under Gaussian).
Given any binary classifier $c:\mathbb{R}^{d}\rightarrow\{-1,+1\}$, a data distribution $\mathcal{N}(0,\mathrm{I})$ and a collection of halfspaces $\mathcal{H}_{1/2}^{\mathcal{N}}$ over $\mathbb{R}^{d}$, there exists an auditing algorithm that takes $\delta>0$ and $N=d^{O(1/\epsilon^{4})}\mathrm{poly}(1/\epsilon)\log(1/\delta)$ labeled i.i.d. example from $\mathcal{N}(0,\mathrm{I})$ in the form of $(\mathbf{x},c(\mathbf{x}))$, runs in time $\mathrm{poly}(N)+(1/\epsilon)^{O(1/\epsilon^{6})}\log(1/\delta)$, and returns a halfspace $h^{\prime}$ as a certificate such that $\Pr_{\mathbf{x}\sim\mathcal{D}}\{h^{\prime}\}=1/2$ and
$\left|d_{\mathcal{N}}(c,h^{\prime})\right|\geq\max_{h\in\mathcal{H}_{1/2}^{\mathcal{N}}}\left|d_{\mathcal{N}}(c,h)\right|-\epsilon$
with at least $1-\delta$ probability.
Proof.
Simply running algorithm 1 with the same set of parameters except that $\mathcal{D}=\mathcal{N}(0,\mathrm{I})$, $n=1$, $a=b=1/2$ and the oracle being as described by lemma 4.2 will give us the desired results. ∎
5 FUTURE WORK
The major drawback of our result is still the lack of approaches of learning halfspaces with sub-constant error guarantee for more general distributions. Therefore, a major direction for fairness auditing remains to develop an agnostic learning method with additive error guarantees for broader classes, such as log-concave distributions—subject to the constraints of Corollary 3.11/[17]. Even a computationally efficient learning algorithm for general halfspaces that can achieve additive error close to $O(1/\sqrt{\log d})$ under Gaussian distributions would be an interesting improvement.
An alternative direction is to seek stronger guarantees for conjunctions on such families of distributions. Conjunctions are more natural in the context of auditing, and their relative lack of expressive power might enable a better guarantee.